# Equivalence and Separation between Heard-Of and Asynchronous Message-Passing Models

### v1

@@ Abstract
We revisit the relationship between two fundamental models of distributed computation: the asynchronous message-passing model with up to $f$ crash failures ($\mathrm{AMP}_{f}$) and the Heard-Of model with up to $f$ message omissions ($\mathrm{HO}_{f}$). We show that for $n>2f$, the two models are equivalent with respect to the solvability of colorless tasks, and that for colored tasks the equivalence holds only when $f=1$ (and $n>2$). The separation for larger $f$ arises from the presence of silenced processes in $\mathrm{HO}_{f}$, which may lead to incompatible decisions. The proofs proceed through bidirectional simulations between $\mathrm{AMP}_{f}$ and $\mathrm{HO}_{f}$ via an intermediate model that captures this notion of silencing. The results extend to randomized protocols against a non-adaptive adversary, indicating that the expressive limits of canonical rounds are structural rather than probabilistic. Together, these results delineate precisely where round-based abstractions capture asynchronous computation, and where they do not.
Keywords: Equivalence of models, Asynchronous Message Passing model, Heard-Of model, Message Adversaries, Task Solvability
@@ 1 Introduction
The standard Asynchronous Message-Passing model with $f$ failures ($\mathrm{AMP}_{f}$) assumes unbounded message delivery times and allows up to $f$ process crashes. Many protocols in this model proceed in rounds: each process broadcasts a message tagged with its current round number and waits for nfn-f messages from that round before advancing to the next. This round-based behavior is formalized in the Heard-Of model [HeardOf], which describes communication through Heard-Of sets, the sets of processes from which each process receives a message in a given round. The variant with $f$ message omissions, denoted $\mathrm{HO}_{f}$, corresponds to executions in which every process hears from at least nfn-f processes per round.
@@
The precise relationship between $\mathrm{HO}_{f}$ and $\mathrm{AMP}_{f}$ has remained unclear. In particular, it is not known whether canonical rounds are fully general, or whether all tasks solvable in $\mathrm{AMP}_{f}$ can be solved also in $\mathrm{HO}_{f}$. Indeed, not every protocol in $\mathrm{AMP}_{f}$ conforms to the canonical-round pattern. For example, the renaming task [attiya_renaming_amp_1990], is solved in $\mathrm{AMP}_{f}$ by a protocol in which each process maintains and shares the set of initial names it knows, and decides once it has received nf1n-f-1 additional copies of this set. Here, decisions depend on the number of identical messages, rather than on the number of messages associated with a particular round. This example highlights that not all asynchronous protocols are expressed in terms of canonical rounds, motivating a closer examination of the precise boundary between $\mathrm{AMP}_{f}$ and $\mathrm{HO}_{f}$.
The most relevant prior work is due to Gafni and Losa [gafni2024timehealer], who showed that $\mathrm{HO}_{1}$ and $\mathrm{AMP}_{1}$ solve the same set of colorless tasks for $n>2$. (Roughly speaking, in a colorless task, like consensus, processes can adopt each other’s decisions.) However, to the best of our knowledge, the case of colored tasks and larger fault thresholds ($f>1$) has not been explored.
@@
We use bidirectional simulations to prove the following results:
@@
$\mathrm{AMP}_{f}$ in $\mathrm{HO}_{f}$ [gafni2024timehealer] Theorem 13 Theorem 7 Theorem 14 unknown
@@
We also show that the results essentially carry over to randomized protocols with private local coins and a non-adaptive adversary, i.e., one whose behavior is fixed independently of the protocol’s random choices; see Theorems 15, 16 and 18. With an adaptive adversary, the separation result still holds because they are stronger than non-adaptive adversaries; however the equivalence results remain open.
@@
A long line of work has sought to bridge asynchronous and round-based message-passing models by placing them within a common framework for reasoning about protocols and impossibility results. The Heard-Of model of Charron-Bost and Schiper [HeardOf], Gafni’s round-by-round fault detectors [gafni_round-by-round_1998], and the unification framework of Herlihy, Rajsbaum, and Tuttle [HerlihyRT98] are central to this direction. The problem of finding a round-based message-passing model equivalent to the read-write wait-free model was addressed by Afek and Gafni [afek2012asynchronysynchrony]. More recently, it was shown that canonical rounds are not fully general, when failures are Byzantine [AttiyaFW25]. Our work contributes to this line by refining the comparison between asynchronous and round-based computation and pinpointing conditions under which equivalence holds.
Another body of research studies how asynchronous message-passing models can be characterized directly in round-based terms. Shimi et al. [ShimiHQ2021] introduce delivered predicates and show how they generate Heard-Of predicates for asynchronous models, providing a formal mechanism for deriving round-based behavior from asynchronous assumptions. Our approach differs in that we focus on task solvability and use the notion of silenced processes to identify where the Heard-Of abstraction $\mathrm{HO}_{f}$ diverges from $\mathrm{AMP}_{f}$.
The Heard-Of abstraction has also proved valuable for the verification of distributed algorithms. Several works, including [ShimiHQ2021, DebratMerz12, BalasubramanianW2020, DamianDMW19, dragoi_logic-based_2014], have used it to reason about correctness within purely safety-based frameworks. Balasubramanian and Walukiewicz [BalasubramanianW2020] formalized this connection, showing that verification in the Heard-Of model can be reduced to reasoning over locally finite trees. Because $\mathrm{HO}_{f}$ is defined by safety predicates and its executions form locally finite trees, automated verification is often more tractable than in $\mathrm{AMP}_{f}$, which also requires liveness properties. The compactness result we establish for $\mathrm{HO}_{f}$ provides a theoretical explanation for this practical advantage.
Beyond verification, other comparisons between message-passing models have focused on the relationship between failure assumptions. Analyses such as Gafni’s round-by-round fault detectors [gafni_round-by-round_1998] investigate reductions between crash and omission failures, highlighting algorithmic reductions between models. Our perspective differs: we study computational equivalence, namely, whether two models can solve exactly the same set of tasks, rather than the existence of specific simulation-based transformations. This shift in focus clarifies the boundaries between asynchronous and round-based computation. In particular, our results refine the Heard-Of framework by identifying when the abstraction of canonical rounds captures the behavior of asynchronous message passing, and when it does not.
@@ 2 Models and preliminaries
We consider message-passing models where processes modeled as deterministic automata with infinite state spaces communicate by broadcasting and receiving messages. A protocol is a set of $n$ processes $\mathcal{P}=\left\{p_{1},\dots,p_{n}\right\}$. We assume our protocols to be full-information, where each process always broadcasts its current view, which is initially its initial state and afterwards its complete local history. A configuration of the system consists of the local states of all the processes and also the state of the environment, e.g., messages in transit, etc. A schedule describes the order in which processes send and receive messages, and which processes’ messages are obtained during each receive. An initial configuration $I$ and a schedule $S$ determine a run $\alpha(I,S)$ of the system which is an alternating sequence of configurations and send or receive events. Each model describes a set of possible schedules and for each schedule, a set of processes that are faulty.
2.1 The $\mathrm{AMP}_{f}$ model
@@
In the Asynchronous Message-Passing Model with $f\geq 0$ process failures ($\mathrm{AMP}_{f}$), there is no fixed upper bound on the time it takes for a message to be delivered nor on the relative speeds of the processors. Processes proceed in atomic steps, where they execute a receive, then a local computation and finally a send. In a send event, denoted as $\operatorname{amp-send}_{i}(m)$, process $p_{i}$ broadcasts message $m$ to all processes including itself. In a receive event, denoted $\operatorname{amp-recv}_{i}(m)$, process $p_{i}$ receives message $m\in M\cup\left\{\bot\right\}$ where $M$ is the message domain and $\bot$ denotes “no message”.
@@
There are multiple definitions for $\mathrm{AMP}$; we present one that is equivalent to the classical definition by Fischer, Lynch and Patterson [FLP] (see Appendix C). A process that has infinitely many steps in a run is said to be non-faulty, and faulty otherwise. There can be at most $f\geq 0$ faulty processes. Messages sent by a non-faulty process are received by all non-faulty processes (Non-faulty Liveness). All messages but the last sent by a faulty process are received by all non-faulty processes (Faulty Quasi-Liveness). Every message received by a process was previously sent to the process (Integrity). No message is received more than once at any process (No Duplicates).
2.2 The $\mathrm{HO}_{f}$ model
@@
The Heard-Of model [HeardOf] is a round-based model. Here we consider one of its instances, the Heard-Of model with $f\geq 0$ message omissions. In each round, every process does a send, then a receive and finally a local computation. In a send event, denoted $\operatorname{ho-send}_{i}(m)$, process $p_{i}$ broadcasts $m$ to all processes including itself. In a receive event, denoted $\operatorname{ho-recv}_{i}(M^{\prime})$, it gets a subset $M^{\prime}$ of the messages sent to it during the round. Messages missed in a round are lost forever.
@@
Conceptually, processes proceed in lockstep; to represent the computation in the form of a sequence, we impose a total order on the events that is consistent with the order of events at each process, e.g., the round-robin pattern [attiyawelch, Section 11.1]. In each round, a process may lose at most $f$ messages from other processes; it always hears from itself (Weak Liveness). Every message received by a process in round $k$ must have been sent in round $k$ by some process (Integrity). A receive event can contain at most one message from each neighbor (No Duplicates). We remark that no process is defined to be faulty.
2.3 Tasks
@@
In a colorless task, processes are free to copy the inputs and outputs of other processes. Formally, let $val(U)$ denote the set of non-$\bot$ values in a vector $U$. In a colorless task, for all input vectors $I,I^{\prime}$ and all output vectors $O,O^{\prime}$ such that $(I,O)\in\Delta$, $val(I)\subseteq val(I^{\prime})$ and $val(O^{\prime})\subseteq val(O)$, we have $(I^{\prime},O)\in\Delta$ and $(I,O^{\prime})\in\Delta$.
@@
A protocol $\mathcal{P}$ solves $T$ in model $\mathcal{M}$, if in every run $R$ of $\mathcal{P}$ with input vector $I$ and output vector $O$, we have (1) $(I,O)\in\Delta$, and (2) $O[i]=\bot$ only if $p_{i}$ is faulty in $\mathcal{M}$ in $R$. Finally, task $T$ is solvable in model $\mathcal{M}$ if there exists a protocol $\mathcal{P}$ that solves $T$ in $\mathcal{M}$.
2.4 Simulations
The definition of simulations follows [attiyawelch, Chapter 7]. A simulation of model $\mathcal{M}_{2}$ in model $\mathcal{M}_{1}$ consists of three layers: (1) $n$ simulated processes $p_{1},\dots,p_{n}$, (2) $n$ simulating machines $P_{1},\dots,P_{n}$, with machine $P_{i}$ assigned to simulated process $p_{i}$, and (3) the communication system of $\mathcal{M}_{1}$ (see Figure 2). Each simulated process $p_{i}$ interacts with its simulating machine $P_{i}$ via $\mathcal{M}_{2}$-communication primitives, as if the communication system is that of $\mathcal{M}_{2}$. The simulating machines interact with each other through $\mathcal{M}_{1}$ using $\mathcal{M}_{1}$-communication primitives. Transitions between states in simulated processes and simulating machines are triggered by the occurrence of events in $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$. The occurrence of an event between $p_{i}$ and $P_{i}$ entails a transition in both $p_{i}$ and $P_{i}$. A simulation is described by specifying the automaton for the simulating machines.
@@ 3 Simulating $\mathrm{HO}_{f}$ in $\mathrm{AMP}_{f}$
We first simulate a variant of the $\mathrm{HO}_{f}$ model, called $\mathrm{CFHO}_{f}$ (for Crashed-Faulty $\mathrm{HO}_{f}$), in $\mathrm{AMP}_{f}$ and then show the inclusion of $\mathrm{HO}_{f}$ in $\mathrm{CFHO}_{f}$. The $\mathrm{CFHO}_{f}$ model is identical to $\mathrm{HO}_{f}$ except that we allow a subset of processes to crash, i.e., have a finite number of events. These processes are considered faulty, so they do not have to decide.
Suppose a protocol $\mathcal{P}$ solves a task $T$ in $\mathrm{CFHO}_{f}$. Each round of $\mathrm{CFHO}_{f}$ can be emulated in $\mathrm{AMP}_{f}$ by waiting for nfn-f messages before proceeding, preserving the causal structure of message delivery. We can construct a simulation system that simulates $\mathrm{CFHO}_{f}$ in $\mathrm{AMP}_{f}$ so that for $1\leq i\leq n$, (1) the $i$-th simulated process is $p_{i}$, and (2) the $i$-th simulation machine $P_{i}$ is specified by Algorithm 1.
@@
2:Initialize empty lists $received[r]$ for all $r\geq 0$
3:When $\operatorname{ho-send}_{i}(m)$ occurs:
@@
5:  Enable $\operatorname{amp-send}_{i}(\left\langle m,round\right\rangle)$
@@
6:When $\operatorname{amp-recv}_{i}(\left\langle m,r\right\rangle)$ occurs:
7:  Add $m$ to $received[r]$
8:Enable $\operatorname{ho-recv}_{i}(received[round])$ when:
9:  $\left\lvert received[round]\right\rvert\geq n-f$
@@
A protocol $\mathcal{Q}$ solving task $T=(\mathcal{I},\mathcal{O},\Delta)$ in $\mathrm{AMP}_{f}$ can be constructed as in Figure 2 by converting the simulation system into a network of automata $\mathcal{Q}=\left\{q_{1},\dots,q_{n}\right\}$ such that $q_{i}$ runs both $p_{i}$ and the $i$-th simulation machine $P_{i}$ internally, and interacts with the communication system of $\mathrm{AMP}_{f}$ using amp-send and amp-recv events. The ho-send and ho-recv events between $p_{i}$ and $P_{i}$ become part of $q_{i}$’s internal computation.
[…]
Figure 1: Simulating $\mathcal{M}_{2}$ in $\mathcal{M}_{1}$
$p_{i}$$P_{i}$$q_{i}$amp-sendho-send(internal)amp-recvho-recv(internal)
Figure 2: Constructing $q_{i}$ from $p_{i}$ and $P_{i}$
@@
The simulator $q_{i}$ works as follows. The input of $q_{i}$ is also used as the input for $p_{i}$. When $p_{i}$ does an internal $\operatorname{ho-send}(m)$ for round $r$ to $P_{i}$, $q_{i}$ does an $\operatorname{amp-send}(\left\langle m,r\right\rangle)$ and waits to receive at least nfn-f messages of the form $\left\langle\cdot,r\right\rangle$, and collects their first components in a set $M^{\prime}$. Then $q_{i}$ performs an internal $\operatorname{ho-recv}(M^{\prime})$ from $P_{i}$ to $p_{i}$ and changes $p_{i}$’s state accordingly. If $p_{i}$ internally decides an output $o$, then $q_{i}$ decides $o$ as well.
@@
In a simulated run in $\mathrm{CFHO}_{f}$, every process $p_{i}$ receives at least nfn-f messages from distinct processes associated with its current round counter before it can increment the counter. This means the simulated run is locally valid at every process, though it need not be globally valid, e.g., a process may finish a simulated round before another starts the same round. However once the local sequence of events are correctly interleaved, the run satisfies the Integrity and No Duplicates properties and so is valid in $\mathrm{CFHO}_{f}$. Thus the vector of outputs (which only depends on the local histories) of all the internal $p_{i}$ and thus all $q_{i}$ is valid for the input vector.
@@
Moreover, it can be shown using induction on the round number that if $q_{i}$ is non-faulty in the run in $\mathrm{AMP}_{f}$, then $p_{i}$ is non-faulty in the simulated run, showing that non-faulty $q_{i}$ gets a $n-\bot o$utput from $p_{i}$. Thus protocol $\mathcal{Q}$ solves task $T$ in $\mathrm{AMP}_{f}$ showing that any task that is solvable in $\mathrm{CFHO}_{f}$ is also solvable in $\mathrm{AMP}_{f}$ for $0\leq f\leq n$.

### v2

@@ Abstract
We revisit the relationship between two fundamental models of distributed computation: the asynchronous message-passing model with up to $f$ crash failures ($\mathrm{AMP}_{f}$) and the Heard-Of model with up to $f$ message omissions ($\mathrm{HO}_{f}$). We show that for $n>2f$, the two models are equivalent with respect to the solvability of colorless tasks, and that for colored tasks the equivalence holds only when $f=1$ (and $n>2$). The separation for larger $f$ arises from the presence of silenced processes in $\mathrm{HO}_{f}$, which may lead to incompatible decisions. The results are proved through bidirectional simulations between $\mathrm{AMP}_{f}$ and $\mathrm{HO}_{f}$, using an intermediate model that captures this notion of silencing. The results extend to randomized protocols against a non-adaptive adversary, indicating that the expressive limits of canonical rounds are structural rather than probabilistic. Together, these results help to delineate where round-based abstractions capture asynchronous computation, and where they do not.
Keywords: Equivalence of models, Asynchronous message-passing model, Heard-Of model, Message adversaries, Task solvability
@@ 1 Introduction
The standard Asynchronous Message-Passing model with $f$ failures ($\mathrm{AMP}_{f}$) assumes unbounded message delivery times and allows up to $f$ process crashes. Many protocols in this model proceed in rounds: each process broadcasts a message tagged with its current round number and waits for nfn-f messages from that round before advancing to the next. This round-based behavior is formalized in the Heard-Of model [10], which describes communication through Heard-Of sets, the sets of processes from which each process receives a message in a given round. The variant with $f$ message omissions, denoted $\mathrm{HO}_{f}$, corresponds to executions in which every process hears from at least nfn-f processes per round.
@@
The precise relationship between $\mathrm{HO}_{f}$ and $\mathrm{AMP}_{f}$ has remained unclear. In particular, it is not known whether canonical rounds are fully general, or whether all tasks solvable in $\mathrm{AMP}_{f}$ can be solved also in $\mathrm{HO}_{f}$. Indeed, not every protocol in $\mathrm{AMP}_{f}$ conforms to the canonical-round pattern. For example, the renaming task [4], is solved in $\mathrm{AMP}_{f}$ by a protocol in which each process maintains and shares the set of initial names it knows and decides once it has received nf1n-f-1 additional copies of this set. Here, decisions depend on the number of identical messages, rather than on the number of messages associated with a particular round. This example highlights that not all asynchronous protocols are expressed in terms of canonical rounds, motivating a closer examination of the precise boundary between $\mathrm{AMP}_{f}$ and $\mathrm{HO}_{f}$.
The most relevant prior work is due to Gafni and Losa [15], who showed that $\mathrm{HO}_{1}$ and $\mathrm{AMP}_{1}$ solve the same set of colorless tasks for $n>2$. (Roughly speaking, in a colorless task, like consensus, processes can adopt each other’s decisions.) However, to the best of our knowledge, the case of colored tasks and larger fault thresholds ($f>1$) has not been explored.
@@
We use bidirectional simulations to prove the following results (see Table 1):
@@
$\mathrm{AMP}_{f}$ in $\mathrm{HO}_{f}$ [15] Theorem 13 Theorem 7 Theorem 14 unknown
@@
We also show that the results carry over to randomized protocols with private local coins and a non-adaptive adversary, i.e., one whose behavior is fixed independently of the protocol’s random choices; see Theorems 15, 16, and 18. With an adaptive adversary, the separation result still holds because they are stronger than non-adaptive adversaries; however, the equivalence results remain open.
@@
A long line of work has sought to bridge asynchronous and round-based message-passing models by placing them within a common framework for reasoning about protocols and impossibility results. The Heard-Of model of Charron-Bost and Schiper [10], Gafni’s round-by-round fault detectors [16], and the unification framework of Herlihy, Rajsbaum, and Tuttle [17] are central to this direction. The problem of finding a round-based message-passing model equivalent to the read-write wait-free model was addressed by Afek and Gafni [1]. More recently, it was shown that canonical rounds are not fully general, when failures are Byzantine [5]. Our work contributes to this line by refining the comparison between asynchronous and round-based computation and pinpointing conditions under which equivalence holds.
Another body of research studies how asynchronous message-passing models can be characterized directly in round-based terms. Shimi, Hurault, and Quéinnec [18] introduce delivered predicates and show how they generate Heard-Of predicates for asynchronous models, providing a formal mechanism for deriving round-based behavior from asynchronous assumptions. Our approach differs in that we focus on task solvability and use the notion of silenced processes to identify where the Heard-Of abstraction $\mathrm{HO}_{f}$ diverges from $\mathrm{AMP}_{f}$.
The Heard-Of abstraction has also proved valuable for the verification of distributed algorithms. Several works, including [18, 12, 7, 11, 13], have used it to reason about correctness within purely safety-based frameworks. Balasubramanian and Walukiewicz [7] formalized this connection, showing that verification in the Heard-Of model can be reduced to reasoning over locally finite trees. Because $\mathrm{HO}_{f}$ is defined by safety predicates and its executions form locally finite trees, automated verification is often more tractable than in $\mathrm{AMP}_{f}$, which also requires liveness properties. The compactness result we establish for $\mathrm{HO}_{f}$ provides a theoretical explanation for this practical advantage.
Beyond verification, other comparisons between message-passing models have focused on the relationship between failure assumptions. Analyses such as Gafni’s round-by-round fault detectors [16] investigate reductions between crash and omission failures, highlighting algorithmic reductions between models. Our perspective differs: we study computational equivalence, namely, whether two models can solve exactly the same set of tasks, rather than the existence of specific simulation-based transformations. This shift in focus clarifies the boundaries between asynchronous and round-based computation. In particular, our results refine the Heard-Of framework by identifying when the abstraction of canonical rounds captures the behavior of asynchronous message passing, and when it does not.
@@ 2 Models and preliminaries
We consider message-passing models where processes modeled as deterministic automata with infinite state spaces communicate by broadcasting and receiving messages. A protocol is a set of $n$ processes $\mathcal{P}=\left\{p_{1},\dots,p_{n}\right\}$. We assume our protocols to be full-information, where each process always broadcasts its current view, which is initially its initial state and afterwards its complete local history. A configuration of the system consists of the local states of all the processes and also the state of the environment, e.g., messages in transit, etc. A schedule describes the order in which processes broadcast and receive messages, and which processes’ messages are obtained during each receive. An initial configuration $I$ and a schedule $S$ determine a run $\alpha(I,S)$ of the system which is an alternating sequence of configurations and broadcast or receive events. Each model describes a set of possible schedules and for each schedule, a set of processes that are faulty.
The $\mathrm{AMP}_{f}$ model
@@
In the Asynchronous Message-Passing Model with $f\geq 0$ process failures ($\mathrm{AMP}_{f}$), there is no fixed upper bound on the time it takes for a message to be delivered nor on the relative speeds of the processors. Processes proceed in atomic steps, where they execute a receive, then a local computation and finally a broadcast. In a broadcast event, denoted as $\operatorname{amp-bc}_{i}(m)$, process $p_{i}$ broadcasts message $m$ to all processes including itself. In a receive event, denoted $\operatorname{amp-recv}_{i}(m)$, process $p_{i}$ receives message $m\in M\cup\left\{\bot\right\}$ where $M$ is the message domain and $\bot$ denotes “no message”.
@@
There are multiple definitions for $\mathrm{AMP}$; we present one that is equivalent to the classical definition by Fischer, Lynch, and Patterson [14] (see Appendix C). A process that has infinitely many steps in a run is said to be non-faulty, and faulty otherwise. There can be at most $f\geq 0$ faulty processes. Messages broadcast by a non-faulty process are received by all non-faulty processes (Non-faulty Liveness). All messages but the last broadcast by a faulty process are received by all non-faulty processes (Faulty Quasi-Liveness). Every message received by a process was previously sent to the process (Integrity). No message is received more than once at any process (No Duplicates).
The $\mathrm{HO}_{f}$ model
@@
The Heard-Of model [10] is a round-based model. Here we consider one of its instances, the Heard-Of model with $f\geq 0$ message omissions. In each round, every process does a broadcast, then a receive, and finally a local computation. In a broadcast event, denoted $\operatorname{ho-bc}_{i}(m)$, process $p_{i}$ broadcasts $m$ to all processes including itself. In a receive event, denoted $\operatorname{ho-recv}_{i}(M^{\prime})$, it gets a subset $M^{\prime}$ of the messages sent to it during the round. Messages missed in a round are lost forever.
@@
Conceptually, processes proceed in lockstep; to represent the computation in the form of a sequence, we impose a total order on the events that is consistent with the order of events at each process, e.g., the round-robin pattern [6, Section 11.1]. In each round, a process may lose at most $f$ messages from other processes; it always hears from itself (Weak Liveness). Every message received by a process in round $k$ must have been broadcast in round $k$ by some process (Integrity). A receive event can contain at most one message from each neighbor (No Duplicates). We remark that no process is defined to be faulty.
Tasks
@@
In a colorless task, processes are free to copy the inputs and outputs of other processes. Formally, let $val(U)$ denote the set of non-$\bot$ values in a vector $U$. In a colorless task, for all input vectors $I,I^{\prime}$ and all output vectors $O,O^{\prime}$ such that $(I,O)\in\Delta$, $val(I)\subseteq val(I^{\prime})$, and $val(O^{\prime})\subseteq val(O)$, we have $(I^{\prime},O)\in\Delta$ and $(I,O^{\prime})\in\Delta$.
@@
A protocol $\mathcal{P}$ solves $T$ in model $\mathcal{M}$, if in every run $R$ of $\mathcal{P}$ with input vector $I$ and output vector $O$, we have (1) $(I,O)\in\Delta$ and (2) $O[i]=\bot$ only if $p_{i}$ is faulty in $\mathcal{M}$ in $R$. Finally, task $T$ is solvable in model $\mathcal{M}$ if there exists a protocol $\mathcal{P}$ that solves $T$ in $\mathcal{M}$.
Simulations
The definition of simulations follows [6, Chapter 7]. A simulation of model $\mathcal{M}_{2}$ in model $\mathcal{M}_{1}$ consists of three layers: (1) $n$ simulated processes $p_{1},\dots,p_{n}$, (2) $n$ simulating machines $P_{1},\dots,P_{n}$, with machine $P_{i}$ assigned to simulated process $p_{i}$, and (3) the communication system of $\mathcal{M}_{1}$ (see Figure 2). Each simulated process $p_{i}$ interacts with its simulating machine $P_{i}$ via $\mathcal{M}_{2}$-communication primitives, as if the communication system is that of $\mathcal{M}_{2}$. The simulating machines interact with each other through $\mathcal{M}_{1}$ using $\mathcal{M}_{1}$-communication primitives. Transitions between states in simulated processes and simulating machines are triggered by the occurrence of events in $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$. The occurrence of an event between $p_{i}$ and $P_{i}$ entails a transition in both $p_{i}$ and $P_{i}$. A simulation is described by specifying the automaton for the simulating machines.
[…]
Figure 1: Simulating $\mathcal{M}_{2}$ in $\mathcal{M}_{1}$
$p_{i}$$P_{i}$$q_{i}$amp-bcho-bc(internal)amp-recvho-recv(internal)
Figure 2: Constructing $q_{i}$ from $p_{i}$ and $P_{i}$
A configuration of the system consists of the states of all the simulated processes and simulating machines and the state of the network. A schedule of a simulation is defined similarly to the schedule of a protocol, except it now involves the primitives of both $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$. A run of a simulation consists of an initial configuration $I$ and a schedule $\alpha$ such that the events in $\alpha$ are enabled in turn starting from $I$.
Given run $\alpha$, we define the simulated run $top(\alpha)$ by restricting the initial configuration of $\alpha$ to the simulated processes and the schedule of $\alpha$ to the events of the upper interface.
@@ 3 Simulating $\mathrm{HO}_{f}$ in $\mathrm{AMP}_{f}$
We first simulate a variant of the $\mathrm{HO}_{f}$ model, called $\mathrm{CFHO}_{f}$ (for Crashed-Faulty $\mathrm{HO}_{f}$), in $\mathrm{AMP}_{f}$ and then show the inclusion of $\mathrm{HO}_{f}$ in $\mathrm{CFHO}_{f}$. The $\mathrm{CFHO}_{f}$ model is identical to $\mathrm{HO}_{f}$ except that we allow a subset of processes to crash, i.e., have a finite number of events. Crashed processes do not need to decide as they are considered faulty processes. The number of messages that a non-crashed process does not receive in a round due to process crashes and message omissions is at most $f$.
Suppose a protocol $\mathcal{P}$ solves a task $T$ in $\mathrm{CFHO}_{f}$. Each round of $\mathrm{CFHO}_{f}$ can be emulated in $\mathrm{AMP}_{f}$ by waiting for nfn-f messages before proceeding, preserving the causal structure of message delivery. We can construct a simulation system that simulates $\mathrm{CFHO}_{f}$ in $\mathrm{AMP}_{f}$ so that for $1\leq i\leq n$, (1) the $i$-th simulated process is $p_{i}$ and (2) the $i$-th simulation machine $P_{i}$ is specified by Algorithm 1.
@@
2:Initialize empty sets $received[r]$ for all $r\geq 0$
3:When $\operatorname{ho-bc}_{i}(m)$ occurs:
@@
5:  Add $m$ to $received[round]$ $\triangleright$ Ensure that $p_{i}$ hears from itself in $\mathrm{HO}_{f}$
6:  Enable $\operatorname{amp-bc}_{i}(\left\langle m,round\right\rangle)$
@@
7:When $\operatorname{amp-recv}_{i}(\left\langle m,r\right\rangle)$ occurs:
8:  Add $m$ to $received[r]$
9:Enable $\operatorname{ho-recv}_{i}(received[round])$ when:
10:  $\left\lvert received[round]\right\rvert\geq n-f$
@@
A protocol $\mathcal{Q}$ solving task $T=(\mathcal{I},\mathcal{O},\Delta)$ in $\mathrm{AMP}_{f}$ can be constructed as in Figure 2 by converting the simulation system into a network of automata $\mathcal{Q}=\left\{q_{1},\dots,q_{n}\right\}$ such that $q_{i}$ runs both $p_{i}$ and the $i$-th simulation machine $P_{i}$ internally, and interacts with the communication system of $\mathrm{AMP}_{f}$ using amp-bc and amp-recv events. The ho-bc and ho-recv events between $p_{i}$ and $P_{i}$ become part of $q_{i}$’s internal computation.
@@
The simulator $q_{i}$ works as follows. The input of $q_{i}$ is also used as the input for $p_{i}$. When $p_{i}$ does an internal $\operatorname{ho-bc}(m)$ for round $r$ to $P_{i}$, $q_{i}$ does an $\operatorname{amp-bc}(\left\langle m,r\right\rangle)$ and waits to receive at least nfn-f messages of the form $\left\langle\cdot,r\right\rangle$, and collects their first components in a set $M^{\prime}$. Then $q_{i}$ performs an internal $\operatorname{ho-recv}(M^{\prime})$ from $P_{i}$ to $p_{i}$ and changes $p_{i}$’s state accordingly. If $p_{i}$ internally decides an output $o$, then $q_{i}$ decides $o$ as well.
@@
Since machine $P_{i}$ waits to receive nfn-f messages tagged with its current round counter before incrementing it, $p_{i}$ receives nfn-f messages in every round of a simulated run before broadcasting in the next round. The Integrity and No Duplicates properties are straightforward to verify and thus the simulated run is locally valid at every process, though it need not be globally valid. Indeed, due to the asynchronous nature of $\mathrm{AMP}_{f}$, machine $P_{i}$ could receive nfn-f round-$r$ tagged messages much earlier than machine $P_{j}$ does, meaning $p_{i}$ could then move on to do a round-$(r+1)$ ho-bc before $p_{j}$ has a ho-recv for round $r$. However once the local history of events at the processes are correctly interleaved to follow the round-robin pattern, the rearranged simulated run is valid in $\mathrm{CFHO}_{f}$. Thus the vector of outputs (which only depends on the local histories) of all the internal $p_{i}$ and thus all $q_{i}$ is valid for the input vector.
We now claim that if $q_{i}$ is non-faulty in a run of $\mathrm{AMP}_{f}$ then $p_{i}$ does not crash in the corresponding simulated run. To see this, let $\mathcal{Q}_{\mathrm{NF}}$ be the set of processes in $\mathcal{Q}$ that are non-faulty in the run of $\mathrm{AMP}_{f}$. Let $q_{i}\in\mathcal{Q}_{\mathrm{NF}}$. In round 0, $q_{i}$ does an internal ho-bc from $p_{i}$ and does a corresponding round-0 tagged amp-bc. As $\left\lvert\mathcal{Q}_{\mathrm{NF}}\right\rvert\geq n-f$, there are nfn-f such amp-bc events. Thus $q_{i}$ receives nfn-f round-0 tagged messages in $\mathrm{AMP}_{f}$ and enables a ho-recv from $P_{i}$ to $p_{i}$; in turn, $p_{i}$ enables a ho-bc for the next round. Thus for every $q_{i}\in\mathcal{Q}_{\mathrm{NF}}$, process $p_{i}$ does a ho-bc for round 1. We can now continue our argument in an inductive fashion to conclude that $p_{i}$ does not crash in the simulated run, which proves the claim.
@@
By the above claim, every non-faulty $q_{i}$ receives a $n-\bot o$utput from $p_{i}$ and hence decides. Thus protocol $\mathcal{Q}$ solves task $T$ in $\mathrm{AMP}_{f}$, showing that any task that is solvable in $\mathrm{CFHO}_{f}$ is also solvable in $\mathrm{AMP}_{f}$ for $0\leq f\leq n$.